OpenStudy (anonymous):

how do work out what x ^2/3 is? (in this case, we're substituting -1 in for x. so -1^2/3 +3

OpenStudy (anonymous):



OpenStudy (anonymous):

x ^(2/3) = (-1)^2/3 = [ cuberoot (-1]) ^2 = (-1) ^2 = 1 1 +3 = 4

OpenStudy (anonymous):

awesome, thank you!

OpenStudy (anonymous):

yw